OpenStudy (anonymous):

20X=28-8Y -8X=-56-8Y

OpenStudy (shinalcantara):

20x = 28 - 8y ------equation 1 -8x = -56 - 8y -----equation 2 ------------------------------ To eliminate y, Subtract equation 2 from equation 1 20x = 28 - 8y - -8x =-56 - 8y 28x = 84 Divide the whole equation with 28 28y/28 = 84/28 x = 3 ---------------------------- Substitute x=3 to either of the two equations above In this case, let's use equation 1 20x = 28 - 8y 20(3) = 28 - 8y 60 = 28 - 8y 60 - 28 = -8y 32 = -8y Divide the whole equation with -8 -4 = y y = -4